Amy wants to fence in a garden whose length is 2m more than twice its width. What is the greatest width her garden can be, if she has at most 70m of fencing?
Answer by ankor@dixie-net.com(12702)

Amy wants to fence in a garden whose length is 2m more than twice its width. What is the greatest width her garden can be, if she has at most 70m of fencing?
:
Let x = the desired width
Then
(2x + 2) = the length
:
2x + 2(2x+2) = 70
2x + 4x + 4 = 70
6x = 70 - 4
6x = 66
x = 66/6
x = 11 m is the desired width
:
Find the length: 2(11) + 2 = 24 m is the length
:
Check solution using the perimeter:
2(24) + 2(11) =
48 + 22 = 70